SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

IFR Remediation 2002

Database for barrel remediation (pdf,xls)
(code definitions, txt)
 
Remediation summaries (.xls files):
Sextant0Sextant1
Sextant2Sextant3
Sextant4Sextant5

SEXTANT 5 (Pre-remediation of August)


Not yet available...

SEXTANT 5 (Official remediation)


Official remediation for Sextant 5 started on Aug.22 at 16:30 (Ar-flow ON) and ended after approx.13 days, on Sep.4 around 12:00.

List of remediated modules:
 Mod.1Mod.2Mod.3
Layer 19POS------
Layer 18NEGPOSPOS
Layer 17---POSNEG
Layer 16---------
Layer 15NEGPOSNEG
Layer 14---------
Layer 13NEGPOSNEG
Layer 12---------
Layer 11---POS---
Layer 10NEGPOS---
Layer 9NEGPOS---
Layer 8POSPOS---
Layer 7NEGPOSNEG
Layer 6---------
Layer 5POS------
Layer 4---POSNEG
Layer 3---------
Layer 2---POS---
Layer 1NEGPOSNEG

Pre-remediation cosmic runs taken on Aug.21st:
Run 297406400 V
Run 297416600 V
Run 297426800 V
Run 297437000 V
Run 297447200 V
Run 297457400 V
Run 297467600 V
Run 297477800 V
Run 297488000 V
Run 297498200 V

Summary of remediation on sextant 5 here.

Plots of current and integrated charge here. Time in plots starts on Aug.22 at 13:00.

Stratigraphic view here.

Post-remediation cosmic runs for efficiency comparison, at 7.6 kV, are:
Run 29925300kevts
Run 299401600kevts!?

SEXTANT 3


Official remediation for Sextant 3 started on Aug.24 at 15:00 (Ar-flow ON) and ended after approx.18 days, on Sep.11 around 12:00.

List of remediated modules:
 Mod.1Mod.2Mod.3
Layer 19---NEGNEG
Layer 18NEGNEGPOS
Layer 17---------
Layer 16---------
Layer 15POSNEGPOS
Layer 14---------
Layer 13---------
Layer 12POSNEG---
Layer 11---------
Layer 10POSNEGPOS
Layer 9POSNEGPOS
Layer 8POSNEG---
Layer 7POS---POS
Layer 6POS------
Layer 5POSNEG---
Layer 4POS---POS
Layer 3---------
Layer 2POSNEG---
Layer 1POSNEG---

Pre-remediation cosmic runs taken on Aug.23rd:
Run 298176400 V
Run 298186600 V
Run 298196800 V
Run 298207000 V
Run 298217200 V
Run 298227400 V
Run 298237600 V
Run 298247800 V
Run 298258000 V
Run 298268200 V

Summary of remediation on sextant 3 here.

Plots of current and integrated charge here. Time in plots starts on Aug.24 at 11:30.

Stratigraphic view here.

Post-remediation cosmic runs at 7600 V only:
Run 29944500kevts
Run 29945500kevts

SEXTANT 2


Official remediation for Sextant 2 started on Sep.24 at 15:30 (Ar-flow ON) and ended after approx.9 days, on Oct.3 around 8:30.

Pre-remediation cosmic run for efficiency comparison is #29986 at 7600V, with 500,000 evts.

Here is the list of remediating RPCs, the currents at nominal HV and mixture before start of remediation, and cartoons for the backward and forward sextant 2, showing past efficiency histories. HV channels PWR1-2-3 (red in the cartoons) were negative, PWR4-5-6 (blue in the cartoons) positive.

Final remediation currents and integrated charges in Sextant 2. Stratigraphic view here. Time in plots starts on Sep.24 at 12:00 (noon).

Summary of remediation on sextant 2 here.

Not all RPC's have integrated as much charge as needed, some for want of Ar flow and some for want of time. Based on the plots, I would recommend to consider for efficiency comparisons only these: 2-7-1, 2-16-1, 2-17-1, 2-18-1, 2-4-2, 2-5-2, 2-12-2, 2-14-2, 2-15-2, 2-16-2, 2-17-2, 2-18-2, 2-11-3, 2-16-3. This is not an official decision, just my feeling.

Post-remediation cosmic run for efficiency comparison is #30167 at 7600V, with ~106,000 evts.

SEXTANT 0


Official remediation for Sextant 0 started on Oct.4 at 10:30 (Ar-flow ON) and ended after approx.10 days, on Oct.14 at 9:00.

Pre-remediation cosmic runs for efficiency comparison are
Run#HV
300707250
300717500
300727750
300738000

RPC's on negative remediation are: 0-4-1, 0-5-1, 0-6-1, 0-7-1, 0-8-1, 0-9-1, 0-10-1, 0-11-1, 0-12-1, 0-14-1, 0-15-1, 0-17-1, 0-18-1, 0-13-2, 0-14-2, 0-15-2, 0-16-2, 0-17-2, 0-18-2

RPC's on positive remediation are: 0-2-2, 0-4-2, 0-6-2, 0-8-2, 0-9-2, 0-10-2, 0-16-1, 0-2-3, 0-6-3, 0-8-3, 0-10-3, 0-13-3, 0-14-3, 0-15-3, 0-16-3, 0-17-3

Remediation for Sextant 0 ended today at 9:00 AM. Final plots for currents and integrated charges are below. (10/14 09:00AM) Remediation currents and integrated charges in Sextant 0. Stratigraphic view here.

Summary of remediation on sextant 0 here.

The Babar gas was restored to Sestant 0 at 2:00 PM and current monitoring has been going on since.

Post-remediation cosmic runs for efficiency comparison are
Run 302546800V40keventstrigg rate 50 Hz.
Run 302557000V100keventstrigg rate 100 Hz.
Run 302567200V100keventstrigg rate 150 Hz.
Run 302577400V100keventstrigg rate 200 Hz.
Run 302587600V100keventstrigg rate 200 Hz.
Run 302597800V100keventstrigg rate 200 Hz.
Run 302608000V100keventstrigg rate 220 Hz.

A second set of after remediation runs was taken on Oct. 18th to compare directly ( runs were taken at the same voltage) to the data collected on Oct. 4th.

Run#HV
303467250
303497500
303507750
303518000
All these runs have 200 Kevents.


SEXTANT 1

Pre-remediation cosmic runs taken on 10/13/2002 for efficiency comparison are
Run#HV
302616800
302627000
302637200
302647400
302657600
302667800
302678000
Sextant 1 gas was changed to pure Argon today (10/14/2002) at 6:30 PM. Modules under remediation are :
Sextant 1
1-1(+) , 1-2(-) , 1-3(+)
2-2(-)
4-1(+)
5-2(-), 5-3(+)
7-1(+),7-2(-),7-3(+)
8-2(-), 8-3(+)
9-2(-)
10-1(+)
12-1(+), 12-2(-)
15-1(-)
16-2(-),16-3(-)
17-3(+)
18-1(-), 18-2(-), 18-3(-)
19-2(+),19-3(+)
Out of these 25 modules 1-17-3 was shorted out, so it was not possible to power it up; 1-19-3 is connected on H.V. distr. box F2 and, as of now, the relative GMB (together with other 3 GMB from CAN01) is not talking. We'll be working to fix this problem so that 24 modules total will be under controlled remediation. Remediation for Sextant 1 ended on 10/24 8:45 AM. After restoring the H.V. and monitoring connection the Sextant was switched to normal gas at 12:15 PM. H.V. was restored around 15:30; the sextant current was too high for the relevant pods before then. The current drop will allow us to measure the flow rate of the (normal) gas into the modules. It is worth noticing that Layers 7,8,13 and 19 that were fed gas both from the input and the output side were NOT restored to this condition, as it was not possible, as of now, to determine from where the gas, fed to the output leg, was coming from.

Remediation currents and integrated charges in Sextant 1 as of the last update below are shown. Stratigraphic view here.

An interesting set of plots can be found here : time history of the individual currents as normal gas is restored to the sextant gives us a better understanding of the gas flow in the sextant layers.

Post-remediation cosmic runs for efficiency comparison are
Run 304626800V
Run 304637000V
Run 304647200V
Run 304657400V
Run 304667600V
Run 304677800V
Run 304688000V


Sextant 4

Pre-remediation cosmic runs taken on 10/19/2002 for efficiency comparison are
Run#HV
303736800
303747000
303757200
303767400
303777600
303787800
303798000
To a first analysis of the reference data, the modules in sextant 4 can be classified as : 0.K. 29, completely dead 8, remediable 20. The situation from the latest data seems better that the one we had at the end of June. Modules that will undergo remediation are : 2-2,3-2,4-1,5-1,5-2,5-3,7-3,8-1,8-2,9-2,9-3,10-1,10-2,13-1,15-1,16-1,18-3,19-1,19-2,19-3.

The breakdown per layer is:
4-2-2 (-)
4-3-2 (-)
4-4-1(-)
4-5-1(-) 4-5-2(-) 4-5-3(+)
4-7-3(+)
4-8-1(-) 4-8-2(-)
4-9-2(-) 4-9-3(+)
4-10-1(-) 4-10-2(-)
4-13-1(-)
4-15-1(-)
4-16-1(-)
4-18-3(+)
4-19-1(+),4-19-2(+),4-19-3(+)

I concentrated the + in the forward and the - in the back. Of the 20 modules I plan to work on, 13 will go negative and 7 positive.

Remediation for Sextant 4 started on 10/24/2002. The new H.V. configuration was implemented at 11:00 AM and currents were recorded with normal gas at 5000V. At 15:00 the sextant was switched to Argon. As for sextant 1 four layers had a double gas feed: 10,13,15 and 19. During Argon flushing the double feed were disconnected.

Remediation currents and integrated charges in Sextant 4 as of the last update below are shown. Stratigraphic view here.


Sextant 4 was switched to normal gas on Nov. 1st at 14:15. Total charge accumulated was on the very low side due to a problem in the gas system and the need to do work on the H.V. system.


Results

A very preliminary analysis of the module efficiencies can be found in here.


Last update on 11/03/2002 by M.Piccolo